OpenStudy (anonymous):

write an equation for the line that is parallel to the given line and that passes through the given point. y=1/2x-8;(-6,-17)

OpenStudy (anonymous):

So you can use the equation Y-y1 = m(X-x1) and to be parallel you need the same slope Y+17=1/2(x-6) Y=1/2x-14